5. Find the i) scalar and ii) vector projections of v onto u if u = 7i+j – 2k and v= 3i -5j +2k
Answer #1

We have a=it j-2K = si-5j+2K is To find s. calar je ū. ✓ =Fit j=2k).(30-55+2K) - 21-5-Y scalar 12 u Liis projection of von au